OpenStudy (raheelafzaal):

∫(sec^2xcosec^2x)dx solve it integration

myininaya (myininaya):

\[\int\limits_{}^{}\sec^2(x) \csc^2(x) dx \] let me think...

myininaya (myininaya):

have you tried integration by parts?

OpenStudy (anonymous):

break one apart?

OpenStudy (raheelafzaal):

yes

OpenStudy (raheelafzaal):

∫(sec2(x)csc2(x))dx

myininaya (myininaya):

Curious, how many rounds did you do?

OpenStudy (raheelafzaal):

what?

OpenStudy (anonymous):

oh wait

OpenStudy (raheelafzaal):

ok

myininaya (myininaya):

omg wait

myininaya (myininaya):

something easier...

OpenStudy (anonymous):

some goofy trig identity that we forget this is \[4\csc^2(2x)\]

OpenStudy (anonymous):

and we know an anti - derivative for \[\csc^2(x)\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{\cos^2(x)} \cdot \frac{1}{\sin^2(x)} dx=\int\limits_{}^{}\frac{1}{(\cos(x) \sin(x))^2} dx\]

myininaya (myininaya):

yeah and you can write whate satellite has :)

OpenStudy (anonymous):

When you get finished could you please look at mine

OpenStudy (raheelafzaal):

thanks

OpenStudy (anonymous):

now mr raheel please look at mine

OpenStudy (raheelafzaal):

ok